OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y −2 −6 −10 −14 y = mx + b m is slope b is y-intercept

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@ganeshie8

OpenStudy (texaschic101):

first find the slope by using slope formula and 2 sets of points slope formula is : slope(m) = (y2 - y1) / (x2 - x1) now use 2 sets of points...it does not matter which ones (1,-2)...x1 = 1 and y1 = -2 (2,-6)...x2 = 2 and y2 = -6 now sub it into the formula slope(m) = (-6-2) / (2-1) slope(m) = -8/1 or just -8 Now we use y = mx + b form and since we already found slope(m) we need to find y intercept(b) y = mx + b slope(m) = -8 use any set of points (1,-2)...x = 1 and y = -2 now sub into the formula and solve for b, the y intercept y = mx + b -2 = -8(1) + b -2 = -8 + b -2 + 8 = b 6 = b so your line in slope intercept form is : y = -8x + 6